OpenStudy (anonymous):

find d^2y/dx^2 for y=(1-x)/(2-x)

OpenStudy (amistre64):

find the second derivative ... right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

try the quotient rule

OpenStudy (amistre64):

then i would either use a product rule with a ^-1; or the quotient rule

OpenStudy (anonymous):

ok thanks so much

OpenStudy (amistre64):

youd have to do it twice tho to get to the second one

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

\[\frac{t}{b}=tb^{-1}\] \[D_x[tb^{-1}]=-tb^{-2}+t'b^{-1}\] \[D_{x}^{2}[tb^{-1}]=D_x[-tb^{-2}+t'b^{-1}]\] \[D_x[-tb^{-2}+b^{-1}]=2tb^{-3}-t'b^{-2}-t'b^{-2}+t''b^{-1}\] witha any luck I kept it all in order lol

OpenStudy (amistre64):

t=(1-x) t' = -1 t'' = 0 b=(2-x) \[2(1-x)(2-x)^{-3}+(2-x)^{-2}+(2-x)^{-2}+0(b^{-1})\] \[2(1-x)(2-x)^{-3}+2(2-x)^{-2}\] \[2(2-x)^{-3}\left((1-x)+(2-x)\right)\] \[\frac{2(3)}{(2-x)^{3}}\] the wolf should be able to check that tho